OpenStudy (anonymous):

solving Special System 3x+y=4 -3x=y-7 Can you guys work it out so i can see how to do it.

OpenStudy (anonymous):

from what i know i think its the substitution method so make y the subject for the first equation: 3x+y=4 y=4-3x then sub in 'y=4-3x' into the 2nd equation -3x=y-7 -3x=(4-3x)-7 find x then substitute in your x into the OTHER equation, not the one you derived it from, else you'll get something like 0=0

OpenStudy (anonymous):

@donmath

OpenStudy (anonymous):

@Green_Plan

OpenStudy (anonymous):

@gudden

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

surd *eye twitches* i think this is an actual question...

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

assuming you typed the equations correctly which i now doubt after attempting to solve >< or maybe its not the substitution method after all oh well

OpenStudy (anonymous):

@Auxuris how do you find x?

OpenStudy (anonymous):

y= 7 - 3x equation one put in the other equation

OpenStudy (anonymous):

well. i don't think its solvable because the xs cancel out each other and i get: 0=-3 or something which certainly can't be right let me try what donmath suggested, though if the first method didn't work, the other way round won't work as well.

OpenStudy (anonymous):

nope i got '7=4' lmao

OpenStudy (gudden):

@surd .. ask a new question n then play the game. this time let the question be " what was the name of first human being "

OpenStudy (anonymous):

yes auxuris i also got 7=4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

omfg people am i missing something here

OpenStudy (anonymous):

that was real funny gudden * laughing hard *

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@gudden

OpenStudy (anonymous):

surd soldier get into your position

OpenStudy (anonymous):

im entering enemys area

OpenStudy (mathmale):

3x+y=4 -3x=y-7 I'd solve this by substitution. Since -3x=y-7, 3x=-(y-7). Substituting the latter result into the first equation: -(y-7) + y = 4 or -y + 7 + y =4 7=4 this is never true, so this system has no solution.

OpenStudy (mathmale):

Several of you respondents were on the right track.